OpenStudy (anonymous):

Solve for x: 3(x + 1) = -2(x - 1) - 4 Answer -1 -25 1 -5

OpenStudy (anonymous):

http://lmgtfy.com/?q=Solve for x: 3(x + 1) = -2(x - 1) - 4

OpenStudy (anonymous):

3x + 3 = -2x + 2 - 4 then 5x = -5 last x = -1